The sum of the measures of two angles is 180 degrees two times the measure of an angle is 21 less than the measure of the angle what is the measure of the smaller angle
Found 3 solutions by Theo, MathTherapy, greenestamps:
Answer by Theo(13342)

the sum of 2 angles is 180 degrees.
2 times the measure of one of the angles is 21 less than the measure of that angle.
what is the measure of the other other angle.
let x equal the larger angle.
let y equal the smaller angle.
x = 2y - 21.
x + y = 180.
relace x with 2y - 21 in the second equation to get:
2y - 21 + y = 180
add 21 to both sides of the equation and combine like terms to get:
3y = 201.
solve for y to get:
y = 201/3 = 67 degrees.
x is equal to 180 minus 67 = 113 degrees.

you get x + y = 67 + 113 = 180 degrees.
you get x = 2y - 21 which becomes 113 = 2*67-21 which becomes 113 = 134-21 which becomes 113 degrees = 113 degrees.
the requirements of the problem are satisfied.
the solution is that the smaller angle is equal to 67 degrees.

Answer by MathTherapy(10552)

The sum of the measures of two angles is 180 degrees two times the measure of an angle is 21 less than the measure of the angle what is the measure of the smaller angle

Answer by greenestamps(13200)



I don't know how the other tutors came up with answers to the problem; the problem as stated is nonsense.

"...two times the measure of an angle is 21 less than the measure of the angle..."

Translation: 2x = x-21 --> x = -21

The measure of an angle (in geometry) can't be negative....

Re-post, stating the conditions correctly.

Note you might find it easier to post the problem correctly if you use some punctuation and sentence structure -- you know, things like periods at ends of sentences and capital letters at the beginning of new sentences.